# Scalable Computations for Generalized Mixed Effects Models with Crossed Random Effects Using Krylov Subspace Methods

### v1

@@ 3 Krylov Subspace Methods for Generalized Mixed Effects Models
Parameter estimation and prediction for models with high-dimensional crossed random effects involve several time-consuming operations. Specifically, there are the following bottlenecks all involving operations with the high-dimensional sparse matrix $\Sigma^{-1}+Z^{T}WZ\in\mathbb{R}^{m\times m}$. First, calculating linear solves $(\Sigma^{-1}+Z^{T}WZ)u=v$, $v\in\mathbb{R}^{m}$, (i) for quadratic forms of log-marginal likelihoods in (3) after applying the Woodbury identity given in (4), (ii) in Newton’s method for finding the mode, see (2), (iii) for implicit derivatives of the log-marginal likelihood, see Proposition 2.1 in Sigrist (2023a), and (iv) for predictive variances given in (12). The latter is particularly challenging as the number of prediction points $n_{p}$ is typically large, and the linear system in (12) thus contains many right-hand sides $Z_{po}^{T}$. Note that predictive variances are required not only for predictive distributions, but also for predictive means of non-Gaussian response variables, see (13). Second, calculating log-determinants $\log\det(\Sigma^{-1}+Z^{T}WZ)$ in (approximate) log-marginal likelihoods $L(y|F,\theta,\xi)$ given in (1) and (5) for non-Gaussian and Gaussian likelihoods, respectively. And third, calculating trace terms such as $\operatorname*{tr}((\Sigma^{-1}+Z^{T}WZ)^{-1}\frac{\partial(\Sigma^{-1}+Z^{T}WZ)}{\partial\theta_{k}})$ for the derivatives of log-determinants and $\operatorname*{tr}(\Psi^{-1}\frac{\partial\Psi}{\partial\theta_{k}}\Psi^{-1}\frac{\partial\Psi}{\partial\theta_{l}})$ for entries of the Fisher information. Traditionally, these operations are performed using a Cholesky decomposition of $\Sigma^{-1}+Z^{T}WZ$. In the following, we show how these operations can be done using Krylov subspace methods.
For linear solves with the matrix $\Sigma^{-1}+Z^{T}WZ$, we use the preconditioned conjugate gradient (CG) method, which solves a linear system $(\Sigma^{-1}+Z^{T}WZ)u=v$ by iteratively doing matrix-vector multiplications with $\Sigma^{-1}+Z^{T}WZ$. This can be done fast since $\Sigma^{-1}+Z^{T}WZ$ is sparse, and convergence is often achieved with $l\ll m$ iterations. Thus, linear solves can be calculated in $O(lm)$ time complexity. For completeness, the preconditioned CG algorithm is included in Appendix A.6.
Several techniques have been proposed in the literature to calculate log-determinants of large, symmetric positive definite matrices. Dong et al. (2017) find that stochastic Lanczos quadrature (SLQ) (Ubaru et al., 2017) achieves the highest accuracy and fastest runtime compared to other methods. In this article, we use the SLQ method in combination with the preconditioned CG method to approximate $\log\det(\Sigma^{-1}+Z^{T}WZ)$. Specifically, we first note that
@@
where $P$ is a symmetric positive definite preconditioner matrix; see Section 3.1 for more information on preconditioners. The last term is approximated with the SLQ method as follows:
@@
where $z_{1},\dots,z_{t}\in\mathbb{R}^{m}$ are i.i.d. random vectors with $\mathbb{E}[z_{i}]=0$ and $\mathbb{E}[z_{i}z_{i}^{T}]=P$, $\tilde{Q}_{i}\tilde{T}_{i}\tilde{Q}_{i}^{T}\approx P^{-\frac{1}{2}}(\Sigma^{-1}+Z^{T}WZ)P^{-\frac{T}{2}}$ is a partial Lanczos decomposition obtained after $l$ steps of the Lanczos algorithm with $P^{-\frac{1}{2}}z_{i}/\|P^{-\frac{1}{2}}z_{i}\|_{2}$ as initial vector, $\tilde{Q}_{i}\in\mathbb{R}^{m\times l}$ has orthonormal columns, $\tilde{T}_{i}\in\mathbb{R}^{l\times l}$ is tridiagonal, and $e_{1}=(1,0,\dots,0)^{T}$. A derivation of this approximation can be found, e.g., in Kündig and Sigrist (2024, Appendix A.2). In this article, we use Gaussian random vectors $z_{i}\sim\mathcal{N}(0,P)$. Wenger et al. (2022) highlight that a decomposition with a preconditioner as in (14) leads to a reduction in variance. Intuitively, the more accurate the preconditioner $P\approx(\Sigma^{-1}+Z^{T}WZ)$, the smaller $\log\det(P^{-\frac{1}{2}}(\Sigma^{-1}+Z^{T}WZ)P^{-\frac{T}{2}})$, and thus the smaller the variance of its stochastic approximation.
As in Gardner et al. (2018), we use a technique from Saad (2003) to calculate the partial Lanczos tridiagonal matrices $\tilde{T}_{1},\dots,\tilde{T}_{t}$ from the coefficients of the preconditioned CG algorithm when solving $(\Sigma^{-1}+Z^{T}WZ)^{-1}z_{1},\dots,(W+\tilde{\Sigma}^{-1})^{-1}z_{t}$ $t$ times; see Appendix A.6. In doing so, we avoid running the Lanczos algorithm, which brings multiple advantages: Numerical instabilities are not an issue, storing $\tilde{Q}_{i}$ is not necessary, and the linear solves $(\Sigma^{-1}+Z^{T}WZ)^{-1}z_{i}$ can be reused in the stochastic trace estimation (STE) for calculating derivatives of the log-determinant, e.g., as follows:
@@
Gradients can thus be calculated with minimal computational overhead once the likelihood is calculated. In Appendix A.4, we show in detail how to calculate derivatives of log-determinants with respect to $\theta$, $F$, $b^{*}$, and $\xi$, using STE and a form of variance reduction with control variates based on the SSOR preconditioner introduced in the next section. Trace terms of the Fisher information in (7) can also be approximated with STE as follows:
$\begin{split}\operatorname*{tr}\left(\Psi^{-1}\frac{\partial\Psi}{\partial\theta_{k}}\Psi^{-1}\frac{\partial\Psi}{\partial\theta_{l}}\right)=&\operatorname*{tr}\left(\left(W-WZ(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}W\right)Z_{k}Z_{k}^{T}\right.\\ &\left.\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \left(W-WZ(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}W\right)Z_{l}Z_{l}^{T}\right)\\ \approx&\frac{1}{t}\sum_{i=1}^{t}\left(\left(Z_{k}Z_{k}^{T}W-Z_{k}Z_{k}^{T}WZ(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}W\right)z_{i}\right)^{T}\\ &\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \left(WZ_{l}Z_{l}^{T}-WZ(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}WZ_{l}Z_{l}^{T}\right)z_{i},\end{split}$ (16)
where we use the Woodbury identity, $\frac{\partial\Psi}{\partial\theta_{k}}=Z_{k}Z_{k}^{T}$, and Gaussian random vectors $z_{i}\sim\mathcal{N}(0,I_{n})$. Linear solves $(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}Wz_{i}$ and $(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}WZ_{l}Z_{l}^{T}z_{i}$ in (16) can be computed with the CG method. In experiments, we have found that an alternative STE for the Fisher information given in Appendix A.5 leads to less accurate estimates than the STE in (16) (results not tabulated).
@@ 3.1 Preconditioners
Preconditioners reduce the variance of approximations of log-determinants and their derivatives and accelerate the convergence speed of the CG method. To practically use a matrix $P$ as a preconditioner, we need to construct it, perform linear solves with it, calculate $\log\det(P)$, and sample from $\mathcal{N}(0,P)$ in a computationally efficient manner. In the following, we describe several preconditioners that we consider in this article.
3.1.1 SSOR preconditioner
The “symmetric successive over-relaxation” (SSOR) preconditioner is given by
@@ 3.1.1 SSOR preconditioner
where $D$ is a diagonal matrix with diagonal entries $(D)_{ii}=(\Sigma^{-1}+Z^{T}WZ)_{ii}$ and $L$ is a strictly lower-triangular matrix with entries $(L)_{ij}=\mathbf{1}_{\{i>j\}}(\Sigma^{-1}+Z^{T}WZ)_{ij}=\mathbf{1}_{\{i>j\}}(Z^{T}WZ)_{ij}$ such that $\Sigma^{-1}+Z^{T}WZ=L+L^{T}+D$.
3.1.2 ZIC preconditioner
The zero fill-in incomplete Cholesky (ZIC) factorization reported in Appendix A.7 gives the ZIC preconditioner $P_{\text{ZIC}}=LL^{T}\approx\Sigma^{-1}+Z^{T}WZ$, where $L\in\mathbb{R}^{m\times m}$ is a sparse lower triangular matrix that has the same sparsity pattern as $\Sigma^{-1}+Z^{T}WZ$. Unfortunately, we sometimes observe breakdowns (Scott and Tůma, 2014), i.e., (clearly) negative numbers in the calculations of square roots.
3.1.3 Diagonal preconditioner
The diagonal preconditioner $P_{\text{Diag}}$ is given by $P_{\text{Diag}}=\text{diag}((\Sigma^{-1}+Z^{T}WZ)_{ii})$. As mentioned in the introduction, this is currently the most popular preconditioner for the CG method for random effects models.
3.1.4 Other preconditioners
We additionally consider two low-rank preconditioners obtained by applying low-rank approximations to the matrix $Z^{T}WZ\approx L_{k}L_{k}^{T}$, $L_{k}\in\mathbb{R}^{m\times k}$ to obtain $P=\Sigma^{-1}+L_{k}L_{k}^{T}$. For these low-rank preconditioners, linear solves and log-determinants can be computed using the Woodbury matrix identity and the matrix determinant lemma, respectively, and the costs thus scale linearly with $m$. First, we apply the pivoted Cholesky decomposition (Harbrecht et al., 2012) to $Z^{T}WZ$ to obtain $L_{k}$ and denote the resulting preconditioner as $P_{\text{PChol}}$. This is a state-of-the-art approach for Gaussian Process regression (Gardner et al., 2018). Second, we apply a partial Lanczos algorithm to approximate $Z^{T}WZ\approx L_{k}L_{k}^{T}$ and denote the corresponding preconditioner as $P_{\text{Lanczos}}$.
@@ 3.2 Predictive variances
The computational bottleneck for prediction is the calculation of the posterior predictive variances $\text{diag}(\Omega_{p})$ in (12). These are required for predictive distributions of the latent and response variables as well as for predictive means of non-Gaussian response variables. Calculating $\text{diag}(\Omega_{p})$ requires solving $n_{p}$ linear equation systems with the matrix $\Sigma^{-1}+Z^{T}WZ$. For high-dimensional crossed random effects and large $n_{p}$, solving these linear systems with the CG method can thus be computationally prohibitive. We develop and compare three solutions for this: two simulation-based methods relying on the CG method and an approach based on the preconditioned Lanczos algorithm.
@@ 3.2.1 Predictive variances using a stochastic estimator for the diagonal of a matrix
Bekas et al. (2007) propose to approximate the diagonal of a matrix $A\in\mathbb{R}^{n_{p}\times n_{p}}$ as $\text{diag}(A)\approx\frac{1}{s}\sum_{i=1}^{s}z_{i}\odot Az_{i}$, where $\odot$ denotes the Hadamard product and $z_{i}\in\mathbb{R}^{n_{p}}$ are Rademacher random vectors with entries ±$\pm 1$. Algorithm 1 presents our method for calculating $\text{diag}(\Omega_{p})$ by stochastically approximating $\text{diag}(Z_{po}(\Sigma^{-1}+Z^{T}WZ)^{-1}Z_{po}^{T})$ in (12) based on the approach of Bekas et al. (2007). Note that Gyger et al. (2024) have used a similar algorithm for predictive covariances in Gaussian process regression. Linear solves for the stochastic estimator are computed with the preconditioned CG method. We additionally apply variance reduction in Algorithm 1 by using a control variate based on $\text{diag}(Z_{po}P^{-1}Z_{po}^{T})$:
@@
where $c\in\mathbb{R}^{n_{p}}$ are the optimal weights for the variance reduction given by $c_{j}=\widehat{\text{Cov}}([h(z_{i})]_{j},[r(z_{i})]_{j})/\widehat{\text{Var}}([r(z_{i})]_{j})$, for $j=1,2,\dots,n_{p}$. This variance reduction is important as otherwise the variance of the stochastic approximation is much larger (results not tabulated). Algorithm 1 can be trivially parallelized and results in an unbiased and consistent approximation for $\text{diag}(\Omega_{p})$; see Appendix A.3 for a proof of Proposition 3.1. Assuming that $P^{-\frac{1}{2}}Z_{po}^{T}$ is precomputed, $P=P^{\frac{1}{2}}P^{\frac{T}{2}}$, the computational complexity of the algorithm is $O(sn_{p}+slm+s\psi)$, where $l$ denotes the number of CG iterations, $s$ the number of simulation iterations, and $\psi$ the number of non-zero entries in $P^{-\frac{1}{2}}Z_{po}^{T}$.
@@
[…]
@@
8:$c_{j}\leftarrow\widehat{\text{Cov}}((z_{i}^{(1)}\odot z_{i}^{(2)})_{j},(z_{i}^{(1)}\odot z_{i}^{(3)})_{j})/\widehat{\text{Var}}((z_{i}^{(1)}\odot z_{i}^{(3)})_{j})$, $j=1,2,\dots,n_{p}$, $c=(c_{1},\dots,c_{n_{p}})^{T}$
9:$\text{diag}(\hat{\Omega}_{p})\leftarrow\text{diag}(Z_{pp}\Sigma_{p}Z_{pp}^{T})+c\odot\text{diag}(Z_{po}P^{-1}Z_{po}^{T})+\frac{1}{s}\sum_{i=1}^{s}(z_{i}^{(1)}\odot z_{i}^{(2)}-c\odot z_{i}^{(1)}\odot z_{i}^{(3)})$
@@
We have also considered a similar algorithm using the alternative version in (11) for $\Omega_{p}$ instead of (12). Specifically, $\text{diag}(Z_{po}\Sigma Z^{T}\Psi^{-1}Z\Sigma Z_{po}^{T})$ is approximated stochastically with the approach of Bekas et al. (2007) using the Woodbury identity, and all other terms are computed deterministically. However, experiments show that such an algorithm results in less accurate estimates compared to Algorithm 1 (results not tabulated).
3.2.2 Predictive covariances using simulation
Adopting an approach proposed in Kündig and Sigrist (2024), predictive covariances $\Omega_{p}$ can be approximated using expression (12) and sampling from a Gaussian distribution with covariance matrix $Z_{po}(\Sigma^{-1}+Z^{T}WZ)^{-1}Z_{po}^{T}$. Algorithm 2 presents this approach. The linear solves $(\Sigma^{-1}+Z^{T}WZ)^{-1}z_{i}^{(3)}$ can be done using the preconditioned CG method and the computational complexity of Algorithm 2 is thus $O(sn+slm+sn_{p})$. This algorithm can also be trivially parallelized and results in an unbiased and consistent approximation for $\Omega_{p}$; see Appendix A.3 for a proof of Proposition 3.2. In contrast to Algorithm 1, this method can be used to calculate the entire covariance matrix and not just the variances on the diagonal. Furthermore, Algorithm 2 can be adapted to compute only the predictive variances $\text{diag}(\boldsymbol{\Omega}_{p})$ by summing $\boldsymbol{z}_{i}^{(4)}\odot\boldsymbol{z}_{i}^{(4)}$ in Line 7. We have also considered applying variance reduction to the estimator in Algorithm 2 by using a control variate based on $Z_{po}P^{-1}Z_{po}^{T}$. However, we found that the variance reduction was very small for predictive variances since $\widehat{\text{Cov}}((z_{i}^{(4)}(z_{i}^{(4)})^{T})_{jk},(z_{i}^{(5)}(z_{i}^{(5)})^{T})_{jk})$, $z_{i}^{(5)}=Z_{po}P^{-\frac{T}{2}}z_{i}^{(1)}$ and $j,k=1,2,\dots,n_{p}$, is often close to zero.

### v2

@@ 3 Krylov Subspace Methods for Generalized Mixed Effects Models
Parameter estimation and prediction for models with high-dimensional crossed random effects require the following time-consuming operations all involving operations with the sparse matrix $\Sigma^{-1}+Z^{T}WZ\in\mathbb{R}^{m\times m}$. First, calculating linear solves $(\Sigma^{-1}+Z^{T}WZ)u=v$, $v\in\mathbb{R}^{m}$, in four tasks: (i) in quadratic forms of log-marginal likelihoods in (3) after applying the Woodbury identity given in (4), (ii) in Newton’s method for finding the mode, see (2), (iii) for implicit derivatives of the log-marginal likelihood, see (sigrist2022latent, Proposition 2.1), and (iv) for predictive variances $\text{diag}(\Omega_{p})$ given in (11). The latter is particularly challenging as the number of prediction points $n_{p}$ is typically large, and $n_{p}$ linear systems need to be solved to calculate $\text{diag}(\Omega_{p})$. Note that predictive variances are required not only for predictive distributions, but also for predictive means of non-Gaussian response variables, see (13). Second, calculating log-determinants $\log\det(\Sigma^{-1}+Z^{T}WZ)$ in (approximate) log-marginal likelihoods $L(y|\beta,\theta,\xi)$ given in (1) and (5) for non-Gaussian and Gaussian likelihoods, respectively. And third, calculating trace terms such as $\operatorname*{tr}((\Sigma^{-1}+Z^{T}WZ)^{-1}\frac{\partial(\Sigma^{-1}+Z^{T}WZ)}{\partial\theta_{k}})$ for the derivatives of log-determinants and $\operatorname*{tr}(\Psi^{-1}\frac{\partial\Psi}{\partial\theta_{k}}\Psi^{-1}\frac{\partial\Psi}{\partial\theta_{l}})$ for entries of the Fisher information. Traditionally, these operations are performed using a Cholesky decomposition of $\Sigma^{-1}+Z^{T}WZ$. In the following, we show how these operations can be done using Krylov subspace methods.
For linear solves with the matrix $\Sigma^{-1}+Z^{T}WZ$, we use the preconditioned conjugate gradient (CG) method, which solves a linear system $(\Sigma^{-1}+Z^{T}WZ)u=v$ by iteratively doing matrix-vector multiplications with $\Sigma^{-1}+Z^{T}WZ$. This can be done fast since $\Sigma^{-1}+Z^{T}WZ$ is sparse, and convergence is often achieved with $l\ll m$ iterations. Thus, linear solves can be calculated in $O(l(m+n))$ time complexity. For completeness, the preconditioned CG algorithm is included in Appendix A.5.
Several techniques have been proposed in the literature to calculate log-determinants of large, symmetric positive definite matrices. dong2017scalable find that stochastic Lanczos quadrature (SLQ) (ubaru2017fast) achieves the highest accuracy and fastest runtime compared to other methods. In this article, we use the SLQ method in combination with the preconditioned CG method to approximate $\log\det(\Sigma^{-1}+Z^{T}WZ)$. Specifically, we first note that
@@
where $P$ is a symmetric positive definite preconditioner matrix; see Section 3.1 for more information on preconditioners. The last term is approximated with the SLQ method (ubaru2017fast) as follows:
@@
where $z_{1},\dots,z_{t}\in\mathbb{R}^{m}$ are i.i.d. random vectors with $\mathbb{E}[z_{i}]=0$ and $\mathbb{E}[z_{i}z_{i}^{T}]=P$, $\tilde{Q}_{i}\tilde{T}_{i}\tilde{Q}_{i}^{T}\approx P^{-\frac{1}{2}}(\Sigma^{-1}+Z^{T}WZ)P^{-\frac{T}{2}}$ is a partial Lanczos decomposition obtained after $l$ steps of the Lanczos algorithm with $P^{-\frac{1}{2}}z_{i}/\|P^{-\frac{1}{2}}z_{i}\|_{2}$ as initial vector, $\tilde{Q}_{i}\in\mathbb{R}^{m\times l}$ has orthonormal columns, $\tilde{T}_{i}\in\mathbb{R}^{l\times l}$ is tridiagonal, and $e_{1}=(1,0,\dots,0)^{T}$. In this article, we use Gaussian random vectors $z_{i}\sim\mathcal{N}(0,P)$. wenger2022preconditioning highlight that a decomposition with a preconditioner as in (14) leads to a reduction in variance. Intuitively, the more accurate the preconditioner $P\approx(\Sigma^{-1}+Z^{T}WZ)$, the smaller $\log\det(P^{-\frac{1}{2}}(\Sigma^{-1}+Z^{T}WZ)P^{-\frac{T}{2}})$, and thus the smaller the variance of its stochastic approximation.
Similarly as in gardner2018gpytorch, we use a technique from saad2003iterative to calculate the partial Lanczos tridiagonal matrices $\tilde{T}_{1},\dots,\tilde{T}_{t}$ from the coefficients of the preconditioned CG algorithm when solving $(\Sigma^{-1}+Z^{T}WZ)^{-1}z_{1},\dots,(\Sigma^{-1}+Z^{T}WZ)^{-1}z_{t}$ $t$ times; see Appendix A.5. In doing so, we avoid running the Lanczos algorithm, which brings multiple advantages: Numerical instabilities of the Lanczos algorithm due to loss of orthogonality (saad2003iterative) are not an issue, storing $\tilde{Q}_{i}$ is not necessary, and the linear solves $(\Sigma^{-1}+Z^{T}WZ)^{-1}z_{i}$ can be reused in stochastic trace estimation (STE) for calculating derivatives of the log-determinant, e.g., as follows:
@@
Gradients can thus be calculated with minimal computational overhead once the likelihood is calculated. In Appendix A.3, we show in detail how to calculate derivatives of log-determinants with respect to $\theta$, $F$, $b^{*}$, and $\xi$, using STE and a form of variance reduction with control variates based on the SSOR preconditioner introduced in the next section.
Trace terms of the Fisher information given in (7) can also be approximated with STE as follows:
$\begin{split}\operatorname*{tr}\left(\Psi^{-1}\frac{\partial\Psi}{\partial\theta_{k}}\Psi^{-1}\frac{\partial\Psi}{\partial\theta_{l}}\right)=&\operatorname*{tr}\left(\left(W-WZ(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}W\right)Z_{k}Z_{k}^{T}\right.\\ &\left.\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \left(W-WZ(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}W\right)Z_{l}Z_{l}^{T}\right)\\ \approx&\frac{1}{t}\sum_{i=1}^{t}\left(\left(Z_{k}Z_{k}^{T}W-Z_{k}Z_{k}^{T}WZ(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}W\right)z_{i}\right)^{T}\\ &\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \left(WZ_{l}Z_{l}^{T}-WZ(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}WZ_{l}Z_{l}^{T}\right)z_{i},\end{split}$ (16)
where we use the Woodbury identity, $\frac{\partial\Psi}{\partial\theta_{k}}=Z_{k}Z_{k}^{T}$, and $z_{i}\sim\mathcal{N}(0,I_{n})$. Linear solves $(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}Wz_{i}$ and $(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}WZ_{l}Z_{l}^{T}z_{i}$ in (16) can be computed with the CG method. In experiments, we have found that an alternative STE for the Fisher information given in Appendix A.4 leads to less accurate estimates than the STE in (16) (results not shown).
@@ 3.1 Preconditioners
Preconditioners reduce the variance of stochastic log-determinant estimators and their derivatives and accelerate the convergence speed of the CG method. The convergence speed and accuracy of the CG and SLQ methods are determined by the spectrum of the preconditioned matrix $P^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P^{-T/2}$; see Section 4 for more details. To practically use a matrix $P$ as a preconditioner, we need to construct it, perform linear solves with it, calculate $\log\det(P)$, and sample from $\mathcal{N}(0,P)$ in a computationally efficient manner. In the following, we describe three preconditioners that we analyze in this article. First, we consider the diagonal preconditioner given by
$P_{\text{Diag}}=\text{diag}((\Sigma^{-1}+Z^{T}WZ)_{ii}).$
Furthemore, the “symmetric successive over-relaxation” (SSOR) preconditioner is given by
@@
where $D$ is a diagonal matrix with diagonal entries $(D)_{ii}=(\Sigma^{-1}+Z^{T}WZ)_{ii}$, and $L$ is a strictly lower-triangular matrix with entries $(L)_{ij}=\mathbf{1}_{\{i>j\}}(\Sigma^{-1}+Z^{T}WZ)_{ij}=\mathbf{1}_{\{i>j\}}(Z^{T}WZ)_{ij}$ such that $\Sigma^{-1}+Z^{T}WZ=L+L^{T}+D$. In addition, the zero fill-in incomplete Cholesky (ZIC) factorization described in Appendix A.6 gives the ZIC preconditioner
$P_{\text{ZIC}}=\tilde{L}\tilde{L}^{T}\approx\Sigma^{-1}+Z^{T}WZ,$
where $\tilde{L}\in\mathbb{R}^{m\times m}$ is a sparse lower triangular matrix that has the same sparsity pattern as $\Sigma^{-1}+Z^{T}WZ$. Unfortunately, we sometimes observe breakdowns (scott2014positive), i.e., (clearly) negative numbers in the calculations of square roots.
In Section 4, we theoretically analyze the SSOR and diagonal preconditioners, and in Section 5, we empirically compare the different preconditioners. We have also considered two low-rank preconditioners $P=\Sigma^{-1}+L_{k}L_{k}^{T}\approx\Sigma^{-1}+Z^{T}WZ$, where $L_{k}\in\mathbb{R}^{m\times k}$ obtained by (i) using the pivoted Cholesky decomposition (harbrecht2012low; gardner2018gpytorch) to approximate $Z^{T}WZ\approx L_{k}L_{k}^{T}$ and (ii) applying a partial Lanczos algorithm to approximate $Z^{T}WZ\approx L_{k}L_{k}^{T}$. However, both these low-rank preconditioners lead to slow CG convergence and high variances of the SLQ method (results not shown).
@@ 3.2 Predictive variances
The computational bottleneck for prediction is the calculation of the posterior predictive variances $\text{diag}(\Omega_{p})$ in (11). These are required for predictive distributions of the latent and response variables as well as for predictive means of non-Gaussian response variables. Calculating $\text{diag}(\Omega_{p})$ requires solving $n_{p}$ linear equation systems with the matrix $\Sigma^{-1}+Z^{T}WZ$. For high-dimensional crossed random effects and large $n_{p}$, solving these linear systems can thus be computationally prohibitive despite the use of the CG method. We develop and compare three solutions for this: two stochastic estimators relying on the CG method and an approach based on the preconditioned Lanczos algorithm.
@@ 3.2.1 Predictive variances using a stochastic estimator for the diagonal of a matrix
Algorithm 1 presents a method for calculating $\text{diag}(\Omega_{p})$ by stochastically approximating $\text{diag}(Z_{po}(\Sigma^{-1}+Z^{T}WZ)^{-1}Z_{po}^{T})$ in (12). The latter is based on an approach of bekas2007estimator for approximating the diagonal of a matrix $A\in\mathbb{R}^{n_{p}\times n_{p}}$ as $\text{diag}(A)\approx\frac{1}{s}\sum_{i=1}^{s}z_{i}\odot Az_{i}$, where $\odot$ denotes the Hadamard product, and $z_{i}\in\mathbb{R}^{n_{p}}$ are Rademacher random vectors with entries $\pm 1$. Linear solves for the stochastic estimator are computed with the preconditioned CG method. We additionally apply variance reduction in Algorithm 1 by using a control variate based on $\text{diag}(Z_{po}P^{-1}Z_{po}^{T})$:
@@
where $c\in\mathbb{R}^{n_{p}}$ are the optimal weights for the variance reduction given by $c_{j}=\widehat{\text{Cov}}([h(z_{i})]_{j},[r(z_{i})]_{j})/\widehat{\text{Var}}([r(z_{i})]_{j})$, for $j=1,2,\dots,n_{p}$. This variance reduction is important for this stochastic estimator, as otherwise the variance of the stochastic approximation is much larger (results not shown). Algorithm 1 can be trivially parallelized and results in an unbiased and consistent approximation for $\text{diag}(\Omega_{p})$; see Appendix A.2 for a proof of Proposition 3.1. Assuming that $P^{-\frac{1}{2}}Z_{po}^{T}$ is precomputed, $P=P^{\frac{1}{2}}P^{\frac{T}{2}}$, the computational complexity of the algorithm is $O(sn_{p}+slm+s\psi)$, where $l$ denotes the number of CG iterations, $s$ the number of simulation iterations, and $\psi$ the number of non-zero entries in $P^{-\frac{1}{2}}Z_{po}^{T}$.
@@
4:  Sample $z_{i}^{(1)}\overset{\text{i.i.d.}}{\sim}\text{Rademacher}$, $z_{i}^{(1)}\in\mathbb{R}^{n_{p}}$
@@
[…]
@@
We have also considered a similar algorithm using the alternative version in (11) for $\Omega_{p}$ instead of (12). Specifically, $\text{diag}(Z_{po}\Sigma Z^{T}\Psi^{-1}Z\Sigma Z_{po}^{T})$ is approximated stochastically with the approach of bekas2007estimator using the Woodbury identity, and all other terms are computed deterministically. However, experiments show that such an algorithm results in less accurate estimates compared to Algorithm 1 (results not shown). Note that gyger2024iterative also use an approach based on bekas2007estimator for predictive covariances in Gaussian process regression.
3.2.2 Predictive covariances by simulating from multivariate Gaussian distributions
Algorithm 2 presents an approach that allows for approximating predictive covariances $\Omega_{p}$ by sampling from a Gaussian distribution with covariance matrix $Z_{po}(\Sigma^{-1}+Z^{T}WZ)^{-1}Z_{po}^{T}$. The linear solves $(\Sigma^{-1}+Z^{T}WZ)^{-1}z_{i}^{(3)}$ can be done using the preconditioned CG method and the computational complexity of Algorithm 2 is thus $O(sn+slm+sn_{p})$. This algorithm can also be trivially parallelized and results in an unbiased and consistent approximation for $\Omega_{p}$; see Appendix A.2 for a proof of Proposition 3.2. In contrast to Algorithm 1, this method can be used to calculate the entire covariance matrix and not just the variances on the diagonal. Furthermore, Algorithm 2 can be adapted to compute only the predictive variances $\text{diag}(\boldsymbol{\Omega}_{p})$ by summing $\boldsymbol{z}_{i}^{(4)}\odot\boldsymbol{z}_{i}^{(4)}$ in Line 7. We have also considered applying variance reduction to the estimator in Algorithm 2 by using a control variate based on $Z_{po}P^{-1}Z_{po}^{T}$. However, we found that the variance reduction was very small for this estimator and control variate since $\widehat{\text{Cov}}((z_{i}^{(4)}(z_{i}^{(4)})^{T})_{jk},(z_{i}^{(5)}(z_{i}^{(5)})^{T})_{jk})$, $z_{i}^{(5)}=Z_{po}P^{-\frac{T}{2}}z_{i}^{(1)}$ and $j,k=1,2,\dots,n_{p}$, is often close to zero. kundig2024iterative has used a similar approach in Gaussian process regression.